OpenStudy (anonymous):

Two in one :) Solve the equation. In (x+5)= In (x-1) - In (x+1) & 400/1+e^-t = 350

OpenStudy (anonymous):

\[\frac{x-1}{x+1}=x+5\]\[(x+1)(x+5)=x-1\]\[x^2+5x+x=x-1\]

myininaya (myininaya):

\[\ln(x+5)=\ln(\frac{x-1}{x+1})\] \[x+5=\frac{x-1}{x+1}\]

OpenStudy (anonymous):

\[x^2+5x+6=0\]

OpenStudy (anonymous):

for the second problem I would start by multiplying the whole equation by 350e^t since e^-t is really 1/e^t

OpenStudy (anonymous):

(x+2)(x+3)=0

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

\[140,000e ^{t}+350=e ^{t}\]

OpenStudy (anonymous):

I do right

OpenStudy (anonymous):

wait a sec the second problem has no real solution to it

OpenStudy (anonymous):

err t=

myininaya (myininaya):

there is a chance he meant to write it as \[\frac{400}{1+e^{-t}}=350\]

OpenStudy (anonymous):

oh right that makes sense

myininaya (myininaya):

400=350(1+e^{-t}) 400=350+350e^{-t} 400e^t=350e^t+350 400e^t-350e^t=350 50e^t=350 e^t=7 and so on...

OpenStudy (anonymous):

t = ln(7)

OpenStudy (anonymous):

\[\frac{x-1}{x+1}=x+5\](x+1)(x+5)=x-1\[x^2+5x+6=0\]\[(x+2)(x+3)=0\] x=-2 x=-3

OpenStudy (anonymous):

i keep trying to solve problems that are written wrong

OpenStudy (anonymous):

in all fairness it wouldnt be logical for 400/1 to appear in a problem

OpenStudy (anonymous):

You guys are awesome !! thanks

OpenStudy (anonymous):

\[ \log_{b}\sqrt[3]{x ^{2\sqrt{y}}}\div z ^{3} \] anyone?? The z^3 is divided under the radical

OpenStudy (anonymous):

this is not a problem. this is an expression

OpenStudy (anonymous):

btw what a weird way to solve \[\frac{400}{1+e^{-t}}=350\]

OpenStudy (anonymous):

ok maybe not